OpenStudy (anonymous):

square root of x + 2 end root + 4 = x

hartnn (hartnn):

like this ? \(\sqrt{x+2}+4=x\) then subtract 4 from both sides, what u get ?

hartnn (hartnn):

only after you have subtracted 4 from both sides, you can square both sides and viola! square roots gets eliminated :)

OpenStudy (anonymous):

but what about the x? I have to get x by itself ! :(

OpenStudy (souvik):

do you subtract 4 from both sides...?

OpenStudy (anonymous):

yeaah so after I subtracted, I got squareroot of x + 2 = x-4

OpenStudy (souvik):

now square both sides..

OpenStudy (anonymous):

x + 2 = x^2 - 8x + 16 ?

OpenStudy (souvik):

get all the terms in one side and try to factor it..

OpenStudy (anonymous):

0 = x^2 - 9x +14

OpenStudy (souvik):

now factor it

OpenStudy (anonymous):

uhmmm, how do I factor this ?

OpenStudy (souvik):

find 2 numbers which adds to -9 and multiplied to 14

OpenStudy (anonymous):

oooh waaait ! (x-7) (x-2)

OpenStudy (souvik):

well done... now....(x-7)(x-2)=0 so (x-7)=0 and (x-2)=0 so you will get x=7 and 2

hartnn (hartnn):

there's more to this problem it has only one solution

OpenStudy (anonymous):

oooh why ?

hartnn (hartnn):

plug in x=7 and x=2 in your original equation and find out by yourself :)

OpenStudy (anonymous):

oooh so the answers just x=7 ! right ?

hartnn (hartnn):

what do you get for left side when x= 2 ?

hartnn (hartnn):

correct! just x= 7 :)

OpenStudy (anonymous):

thank you both !

hartnn (hartnn):

welcome dear :)